The wavelength associated with a golf ball weighing 200g and moving at a speed of 5m/h is of the order

Grade:11

1 Answers

Anirudh Natarajan
6 years ago
De broglie wavelength = h/mvh= 6.626×10^-34m=2×10^-1v=5÷3600= 5×10^-2 ÷366.626×10^-34×36÷2×10^-1×5×10-2Order = 10^xOrder = 10^-30